# Bayesian changepoint detection via logistic regression and the topological analysis of image series

### v2

@@ Abstract
Keywords: Changepoint Analysis; Multivariate Time Series; Nonparametric; Quasi-Bayesian; Persistent Homology
@@ 1 Introduction
The model for the data. Conventional likelihood-based approaches to changepoint analysis require the specification of a model for the data within each homogeneous segment of the observed time series. There is a tradeoff between fidelity to the data and parsimony of the model (which is typically closely related to its computational tractability). On one end of the spectrum, there are simple parametric methods that make restrictive assumptions regarding the distribution of the data within each segment, e.g. that the data are Gaussian (Srivastava and Worsley, 1986; Lavielle and Teyssiere, 2006) or follow an exponential family distribution (Chen and Gupta, 2012). On the other end, there are elaborate Bayesian nonparametric methods that avoid restrictive assumptions but may be difficult to for the non-expert to implement or interpret (Martínez and Mena, 2014; Corradin et al., 2022). Negotiating this tradeoff between fidelity and parsimony becomes much more difficult in the context of multivariate time series, and even more so when the time series include both continuous and discrete components.
@@
To address these challenges, we introduce a new method for Bayesian changepoint analysis in the offline, single changepoint setting. The method allows for simultaneous inference on the location of a changepoint and the coefficients of a logistic regression model for distinguishing pre-changepoint data from post-changepoint data. The regression coefficients provide an interpretable description of a potentially complex change. Because the observed time series is treated as a sequence of covariate vectors, there is no need to specify a model for the data, and the method can be applied to data of mixed type. For posterior inference, the model admits a simple Gibbs sampling algorithm based on Pólya-gamma data augmentation (Polson et al., 2013). We establish conditions under which the proposed method is guaranteed to recover the true underlying changepoint.
@@
As a testing ground for our method, we consider the problem of detecting topological changes in time series of images. Most methods in the image change detection literature consider pixelwise differences or small sequences of images (Radke et al., 2005). These methods fail when faced with substantial noise. One can develop more robust methods by focusing on those quantitative summaries or features of an image series most relevant to detecting a change. Topological data analysis (TDA) has gained traction in the statistics and machine learning communities by developing features that lead to improved classification (Hensel et al., 2021). For example, Turkes et al. (2022) and Obayashi et al. (2018) demonstrate that TDA, and persistent homology in particular, is effective for learning nonlinear features of data in an off-the-shelf fashion. We demonstrate that the proposed changepoint method, in conjunction with a topological feature embedding, successfully recovers the location and topological nature of nonlinear changes in image series.
We now outline the remainder of the article. In Section 2, we introduce the proposed changepoint model and provide theoretical results that shed light on its efficacy. Section 3 reviews important concepts from topological data analysis and describes the feature embedding we use for detecting topological changes in image series. In Section 4, we evaluate the proposed model on simulated and real image data. We find that the performance of our method is comparable to or better than that of the state-of-the-art methods and that our method provides useful information not available from competing methods. Section 5 illustrates the performance of the method in two important changepoint settings. The first involves data of mixed type, while the second involves a change in covariance. The conclusions of Section 5 are consistent with those of Section 4. We conclude in Section 6 with a summary of our main contributions and a discussion of future directions.
@@ 2 Bayesian changepoint detection via logistic regression
We begin, in the first subsection, by introducing the proposed changepoint method. In the second subsection, we present two important theoretical properties that help illuminate how the method works.
@@ 2.1 The changepoint model
The proposed changepoint method can be understood from two complementary perspectives: a Bayesian perspective and a quasi-Bayesian perspective. The Bayesian formulation is nonstandard because the posterior results from conditioning on an event other than the usual observation of data. This nonstandard formulation retains the advantages of a Bayesian approach without requiring us to specify a model for the data within each homogeneous segment of the observed time series. In this sense, our method has something in common with Hoff (2007), Miller and Dunson (2019), and Lewis et al. (2021), each of which propose to overcome some challenge associated with a conventional Bayesian analysis by conditioning on a carefully chosen event. The quasi-Bayesian formulation shares these same advantages but offers more flexibility to incorporate prior information regarding the changepoint. Quasi-Bayesian methods, also known as generalized Bayesian methods (Bissiri et al., 2016), are well-established in the statistics and econometrics literatures, and are typically motivated as a robust, nonparametric alternatives to classical or Bayesian estimation. In the econometrics literature, quasi-Bayesian methods are also referred to as Laplace-type methods (Chernozhukov and Hong, 2003; Casini and Perron, 2022).
@@ 2.1.1 The Bayesian formulation
where $\bm{\beta}\in\mathbbm{R}^{d}$ is a vector of unknown regression coefficients. Taking a Bayesian perspective, we assign $\bm{\beta}$ a Gaussian prior, $\bm{\beta}\sim\text{N}\left(\bm{\mu},\Sigma\right).$ Letting $\bm{Y}=(Y_{1},\ldots,Y_{n})^{\top}$ and $\bm{X}$ be the $n\times d$ matrix whose $i$th row is $\bm{x}_{i}^{\top},$ the prior distribution for $\bm{\beta}$ and the conditional distribution specified in (1) define a joint distribution for $\left(\bm{Y},\bm{\beta}\right)$ with a density $p(\bm{y},\bm{\beta}\mid\bm{X}).$ In the standard logistic regression setting, we would condition on the precise value of the response vector $\bm{Y}=\bm{y}$ to get the posterior density $p\left(\bm{\beta}\mid\bm{Y}=\bm{y},\bm{X}\right).$ Tuning-free posterior inference could then be carried out with Pólya-Gamma data augmentation as described in Polson et al. (2013).
@@
=$=\left\{\prod_{i=1}^{\kappa}\frac{1}{1+e^{\bm{x}_{i}^{\top}\bm{\beta}}}\prod_{i=\kappa+1}^{n}\frac{e^{\bm{x}_{i}^{\top}\bm{\beta}}}{1+e^{\bm{x}_{i}^{\top}\bm{\beta}}}\right\}\,\pi_{\text{B}}\left(\bm{\beta}\right)$
@@
We can derive a Gibbs sampler for posterior simulation by adapting the Pólya-Gamma data augmentation scheme of Polson et al. (2013). We augment the parameter space with a vector $\bm{\omega}=\left(\omega_{1},\ldots,\omega_{n}\right)$ of Pólya-Gamma latent variables, yielding a posterior density over the parameters $\left(\kappa,\bm{\beta},\bm{\omega}\right)$ with tractable full conditional distributions. The full conditional distributions for $\bm{\beta}$ and the $\omega_{i}$’s are identical to those appearing in Polson et al. (2013). The full conditional distribution for $\kappa$ is a discrete distribution supported on the set $\{1,\ldots,n-1\}$ and satisfies
@@ 2.1.2 The quasi-Bayesian formulation
The proposed changepoint method can be also be understood as quasi-Bayesian in the sense of Chernozhukov and Hong (2003). In the quasi-Bayesian formulation, we directly define a quasi-likelihood
@@
$\pi_{\text{QB}}(\kappa,\bm{\beta}\mid\bm{X})\propto\mathcal{Q}(\kappa,\bm{\beta}\mid\bm{X})\pi_{\text{QB}}(\kappa,\bm{\beta}).$
The quasi-likelihood is bounded above by one, implying that a proper prior yields a proper quasi-posterior. If we suppose that $\pi_{\text{QB}}(\kappa,\bm{\beta})=\pi_{\text{QB}}\left(\kappa\right)\pi_{\text{QB}}\left(\bm{\beta}\right),$ we can simulate from the from the quasi-posterior with a Gibbs sampler nearly identical to that of the previous section. The full conditional distribution for $\kappa$ becomes
@@ 3 Topological analysis of image data
In this section, we provide a very cursory introduction to the particular TDA method we use in this paper to derive topological information: the persistent homology of images. We then discuss how we go about using the derived persistence diagrams to choose a vectorization (i.e. multivariate feature representation) which captures topological changes in image series.
@@ 3.1 Persistent homology of images
Often it is convenient to treat a ($k\times l$) image as a vector $\tilde{\bm{x}}_{i}$ in $\mathbb{R}^{kl}$. However, for the computation of shape information via persistent homology, it is more convenient to treat $\tilde{\bm{x}}_{i}$ as a function with finite, rectangular support from the two-dimensional integer lattice $\mathbb{Z}^{2}$ to the extended reals $\bar{\mathbb{R}}$—see Section S5 in the Supplementary Material for more details. The sublevel sets of such a function can be considered as binary images, and cubical homology can be calculated from them (cf. Chapter 2, Kaczynski et al., 2004). Keeping track of the cubical homology of each successive binary image and how the connected components merge, and holes fill in, yields the persistence diagrams $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$ respectively [^1] The persistence diagrams we calculate from our images contain all of the information associated to the cubical persistent homology of said images. . Persistence diagrams $\mathcal{D}$ in this context are multisets of points $(b,d)$ in $\mathbb{R}^{2}$ where the $x$ and $y$ coordinates of $(b,d)\in\mathcal{D}$ correspond to the pixel values at which a shape feature appears (is born) and then merges/fills in (dies)—see Figure S4 for an illustration.
Remark 3.1.
In a binary image, a connected component (contributing to $\mathcal{D}^{0}$) is a connected black region and a loop/hole (contributing to $\mathcal{D}^{1}$) is a white region surrounded by black pixels. Note that for the purpose of computing homology we consider pixels outside of a binary image as white.
More information on the finer points of how we compute persistent homology and precisely how images are dealt with appears in Section S5 of the Supplementary Material. Excellent surveys on TDA for statistical/machine learning purposes can be seen in Chazal and Michel (2021) and Hensel et al. (2021).
@@ 3.3 Crafting a topological feature mapping
The approach that we consider is to select a topological feature embedding (from Chung et al. 2018) which represents a wide range of topological summary statistics, wherein the statistics we describe were shown to achieve strong test accuracy in a support vector machine classification task of skin lesions. We will use the same suite of persistence statistics used by Chung et al. (2018), along with the ALPS statistic introduced in Thomas et al. (2023). Namely, we define $l_{i}=d_{i}-b_{i}$ and $m_{i}=(d_{i}+b_{i})/2$ and construct our topological embedding $f_{\text{stat}}$ to have the form
@@
with $T_{j}$ equal to the following statistics of the empirical distributions of $l_{i}$ and $m_{i}$ for the persistence diagrams $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$. The various $T_{j}$ are the mean of $l_{i}$ and $m_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; variance of $l_{i}$ and $m_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; skewness of $l_{i}$ and $m_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; kurtosis of $l_{i}$ and $m_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; $\bm{25^{th}}$, $\bm{50^{th}}$, and $\bm{75^{th}}$ percentiles of $l_{i}$ and $m_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; interquartile range of $l_{i}$ and $m_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; persistent entropy of $l_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; and ALPS statistic of $l_{i}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$.
@@
4 Simulation & applications: TDA and images
To demonstrate the utility of our method, we consider a simple simulated setup and evaluate our method against other well-known methods in the multivariate changepoint literature. We consider both a straightforward and a more difficult (noisier) changepoint problem for a topological change in a rather short image series. This is to show that our topological feature embedding needs relatively little data to perform well, and that our changepoint method also performs well and is robust in this “small” data setting.
4.1 Simulation

### v3

@@ Abstract
Keywords: Changepoint Analysis; Multivariate Time Series; Nonparametric; Quasi-Bayesian; Persistent Homology; Generalized Bayes
@@ 1 Introduction
The model for the data. Conventional likelihood-based approaches to changepoint analysis require the specification of a model for the data within each homogeneous segment of the observed time series. There is a tradeoff between fidelity to the data and parsimony of the model (which is typically closely related to its computational tractability). On one end of the spectrum, there are simple parametric methods that make restrictive assumptions regarding the distribution of the data within each segment, e.g. that the data are Gaussian (Srivastava and Worsley, 1986; Lavielle and Teyssiere, 2006) or follow an exponential family distribution (Chen and Gupta, 2012). On the other end, there are elaborate Bayesian nonparametric methods that avoid restrictive assumptions but may be difficult for the non-expert to implement or interpret (Martínez and Mena, 2014; Corradin et al., 2022). Negotiating this tradeoff between fidelity and parsimony becomes much more difficult in the context of multivariate time series, and even more so when the time series include both continuous and discrete components.
@@
To address these challenges, we introduce a new method for Bayesian changepoint analysis in the offline setting called BCLR. Though the method we devise is tailored to the setting where there is at most one changepoint, we detail an extension to the multiple changepoint setting in Section 6 and present results of the multiple changepoint method in Section S6 of the Supplementary Material. Our method allows for simultaneous inference on the location of a changepoint and the coefficients of a logistic regression model for distinguishing pre-changepoint data from post-changepoint data. The regression coefficients provide an interpretable description of a potentially complex change. Because the observed time series is treated as a sequence of covariate vectors, there is no need to specify a model for the data, and the method can be applied to data of mixed type. For posterior inference, the model admits a simple Gibbs sampling algorithm based on Pólya-gamma data augmentation (Polson et al., 2013). We establish conditions under which the proposed method is guaranteed to recover the true underlying changepoint.
@@
As a testing ground for our method, we consider the problem of detecting topological changes in time series of images. Most methods in the image change detection literature consider pixelwise differences or small sequences of images (Radke et al., 2005). These methods fail when faced with substantial noise. One can develop more robust methods by focusing on those quantitative summaries or features of an image series most relevant to detecting a change. Topological data analysis (TDA) has gained traction in the statistics and machine learning communities by developing features that lead to improved classification (Hensel et al., 2021). For example, Turkes et al. (2022) and Obayashi et al. (2018) demonstrate that TDA, and persistent homology in particular, is effective for learning various nonlinear features of image data in an off-the-shelf fashion. We demonstrate that the proposed changepoint method, in conjunction with a topological feature embedding, successfully recovers the location and topological nature of nonlinear changes in image series.
We now outline the remainder of the article. In Section 2, we introduce the proposed changepoint model and provide theoretical results that shed light on its efficacy. Section 3 reviews important concepts from topological data analysis and describes the feature embedding we use for detecting topological changes in image series. In Section 4, we evaluate the proposed model on simulated image data and in two other important changepoint settings. The first involves data of mixed type, while the second involves a change in covariance. We find that the performance of our method is comparable to or better than that of the state-of-the-art methods and that our method provides useful information not available from competing methods. In Section 5, we evaluate the method on real-world image data, and the conclusions we see are consistent with those of Section 4. Section 6 details an extension to the multiple changepoint setting which retains the fundamental features of the single changepoint method as well as its competitive performance. We conclude in Section 7 with a summary of our main contributions and a discussion of future directions.
@@ 2 Bayesian changepoint detection via logistic regression
We begin, in the first subsection, by introducing the proposed changepoint method for the single changepoint setting. In the second subsection, we present two important theoretical properties that help illuminate how the method works.
@@ 2.1 The changepoint model
The proposed changepoint method can be understood from two complementary perspectives: a Bayesian perspective and a quasi-Bayesian perspective. The Bayesian formulation is nonstandard because the posterior results from conditioning on an event other than the usual observation of data. This nonstandard formulation retains the advantages of a Bayesian approach without requiring us to specify a model for the data within each homogeneous segment of the observed time series. In this sense, our method has something in common with Hoff (2007), Miller and Dunson (2019), and Lewis et al. (2021), each of which propose to overcome some challenge associated with a conventional Bayesian analysis by conditioning on a carefully chosen event. The quasi-Bayesian formulation shares these same advantages but offers more flexibility to incorporate prior information regarding the changepoint. Quasi-Bayesian methods, also known as generalized Bayesian methods (Bissiri et al., 2016), are well-established in the statistics and econometrics literatures, and are typically motivated as a robust, nonparametric alternative to classical or Bayesian estimation. In the econometrics literature, quasi-Bayesian methods are also referred to as Laplace-type methods (Chernozhukov and Hong, 2003; Casini and Perron, 2022).
@@ 2.1.1 The Bayesian formulation
where $\bm{\beta}\in\mathbbm{R}^{d}$ is a vector of unknown regression coefficients. Taking a Bayesian perspective, we assign $\bm{\beta}$ a Gaussian prior, $\bm{\beta}\sim\text{N}\left(\bm{\mu},\Sigma\right).$ Letting $\bm{Y}=(Y_{1},\ldots,Y_{n})^{\top}$ and $\bm{X}$ be the $n\times d$ matrix whose $i^{\text{th}}$ row is $\bm{x}_{i}^{\top},$ the prior distribution for $\bm{\beta}$ and the conditional distribution specified in (1) define a joint distribution for $\left(\bm{Y},\bm{\beta}\right)$ with a density $p(\bm{y},\bm{\beta}\mid\bm{X}).$ In the standard logistic regression setting, we would condition on the precise value of the response vector $\bm{Y}=\bm{y}$ to get the posterior density $p\left(\bm{\beta}\mid\bm{Y}=\bm{y},\bm{X}\right).$ Tuning-free posterior inference could then be carried out with Pólya-Gamma data augmentation as described in Polson et al. (2013).
@@
=$=\left\{\prod_{i=1}^{\kappa}\frac{1}{1+e^{\bm{x}_{i}^{\top}\bm{\beta}}}\prod_{i=\kappa+1}^{n}\frac{e^{\bm{x}_{i}^{\top}\bm{\beta}}}{1+e^{\bm{x}_{i}^{\top}\bm{\beta}}}\right\}\,\pi_{\text{B}}\left(\bm{\beta}\right)$ (2)
@@
We can derive a Gibbs sampler for posterior simulation by adapting the Pólya-Gamma data augmentation scheme of Polson et al. (2013). Data augmentation is necessary because there is no efficient way to directly simulate from the full conditional density $\pi_{\text{B}}\left(\bm{\beta}\mid\kappa,\bm{X}\right)$ arising from (2.1.1). The idea of data augmentation is to augment the parameter space by introducing additional latent variables such that 1) we recover the original posterior distribution when we marginalize over the latent variables and 2) we can easily simulate from the resulting full conditional distributions. The full conditional density $\pi_{\text{B}}\left(\bm{\beta}\mid\kappa,\bm{X}\right)$ is identical to the posterior density of the regression coefficients in a logistic regression model with a multivariate normal prior. Thus, we can leverage Pólya-Gamma data augmentation, which has become a standard approach to posterior simulation in Bayesian methods related to logistic regression. We augment the parameter space with a vector $\bm{\omega}=\left(\omega_{1},\ldots,\omega_{n}\right)$ of Pólya-Gamma latent variables, yielding a posterior density over the parameters $\left(\kappa,\bm{\beta},\bm{\omega}\right)$ with tractable full conditional distributions. The full conditional distributions for $\bm{\beta}$ and the $\omega_{i}$’s are identical to those appearing in Polson et al. (2013). The full conditional distribution for $\kappa$ is a discrete distribution supported on the set $\{1,\ldots,n-1\}$ and satisfies
@@ 2.1.2 The quasi-Bayesian formulation
The proposed changepoint method can also be understood as quasi-Bayesian in the sense of Chernozhukov and Hong (2003). In the quasi-Bayesian formulation, we directly define a quasi-likelihood
@@
$\pi_{\text{QB}}(\kappa,\bm{\beta}\mid\bm{X})\propto\mathcal{Q}(\bm{\beta},\kappa\mid\bm{X})\pi_{\text{QB}}(\kappa,\bm{\beta}).$
The quasi-likelihood is bounded above by one, implying that a proper prior yields a proper quasi-posterior. If we suppose that $\pi_{\text{QB}}(\kappa,\bm{\beta})=\pi_{\text{QB}}\left(\kappa\right)\pi_{\text{QB}}\left(\bm{\beta}\right),$ we can simulate from the quasi-posterior with a Gibbs sampler nearly identical to that of the previous section. The full conditional distribution for $\kappa$ becomes
@@ 3 Topological analysis of image data
Now we provide a very cursory introduction to the particular TDA method we use in this paper to derive topological information: the persistent homology of images. We then discuss how we go about using the derived persistence diagrams to choose a vectorization (i.e. multivariate feature representation) which captures topological changes in image series.
@@ 3.1 Persistent homology of images
Often it is convenient to treat a ($k\times l$) image as a vector $\tilde{\bm{x}}_{i}$ in $\mathbb{R}^{kl}$. However, for the computation of shape information via persistent homology, it is more convenient to treat $\tilde{\bm{x}}_{i}$ as a function—called the image map $I$—with finite, rectangular support on the two-dimensional integer lattice $\mathbb{Z}^{2}$ which takes values in the extended reals $\bar{\mathbb{R}}$. The sublevel sets of such a function can be considered as binary images, and keeping track of the how the shape information in these images changes through thresholding yields the persistence diagrams $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$, corresponding to 0- and 1-dimensional features respectively. More information on the finer points of persistence diagrams and its underlying theory (persistent homology), appear in Section S5 of the Supplementary Material. To continue introducing persistence diagrams, we need to define various shape features in a binary image.
Definition 3.1.
In a binary image, a connected component (contributing to $\mathcal{D}^{0}$) is a connected black region and a loop/hole (contributing to $\mathcal{D}^{1}$) is a connected white region surrounded by black pixels. Note that for the purpose of computing homology we consider pixels outside of a binary image as white.
In general, a persistence diagram $\mathcal{D}$ is a multiset of points $(b,d)$ in $\mathbb{R}^{2}$ where the $x$ and $y$ coordinates of $(b,d)\in\mathcal{D}$ correspond to the pixel values at which a shape feature appears (is born) and then merges/fills in (dies)—see Figure S4 for an illustration. To explain this further, we will take $\mathcal{D}^{0}$ as an example [^1] As the persistence diagram $\mathcal{D}^{0}$, and $\mathcal{D}^{1}$ for the negated image, satisfy a duality property (Garin et al., 2020), we only describe the persistence diagram $\mathcal{D}^{0}$.. For simplicity, we assume that our image map $I$ is injective so that all pixels have a unique intensity value. If the pixel $p$ is such that $I(p)$ is a local minimum, then it will appear at the threshold $b_{p}$ and not be surrounded by any other black pixels. The pixel $p$ creates or “gives birth” to a connected component $C_{p}$. As the threshold $t$ increases, $C_{p}$ will continue to gain more and more black pixels until it merges with another connected component $C_{q}$ at threshold $d$. If the threshold $b_{q}$ at which $C_{q}$ appears is less than $b_{p}$, then we say that $C_{p}$ “dies” at threshold $d_{p}=d$, and associate the values $(b_{p},d_{p})$ to the connected component $C_{p}$. Thus, for the set $M$ of all the local minima in the image we have that
$\mathcal{D}^{0}=\big{\{}(b_{p},d_{p}):p\in M\big{\}},$
is the 0-dimensional persistence diagram for $I$. These can be transformed/vectorized in various ways that make them highly useful in statistical/machine learning contexts; one should consult the excellent surveys by Chazal and Michel (2021) and Hensel et al. (2021) for further examples beyond those given below.
@@ 3.3 Crafting a topological feature mapping
The approach that we consider is to select a topological feature embedding (from Chung et al. 2018) which represents a wide range of topological summary statistics, wherein the statistics we describe were shown to achieve strong test accuracy in a support vector machine classification task of skin lesions. We will use the same suite of persistence statistics used by Chung et al. (2018), along with the ALPS statistic introduced in Thomas et al. (2023). To further justify our choice in this context, we note that the ALPS statistic and the persistent entropy were demonstrated to capture nanoparticle dynamics well in the videos of Thomas et al. (2023). Additionally, the same slate of persistence statistics we use here demonstrated superior classification ability in the microscopy imaging study of Pritchard et al. (2023). Furthermore, as the persistence diagrams $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$ provide information on the connectivity structure of dark (resp. light) regions in the grayscale image series we analyze, these persistence statistics provide us interpretable information on the distribution of this connectivity.
To detail the persistence statistics we used, first define $l_{p}=d_{p}-b_{p}$ and $m_{p}=(d_{p}+b_{p})/2$, $p\in M$ and construct our topological embedding $f_{\text{stat}}$ to have the form
@@
with $T_{j}$ equal to the following statistics of the empirical distributions of $l_{p}$ and $m_{p}$ for the persistence diagrams $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$. The various $T_{j}$ are the means of $l_{p}$ and $m_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; variance of $l_{p}$ and $m_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; skewness of $l_{p}$ and $m_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; kurtosis of $l_{p}$ and $m_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; $25^{th}$, $50^{th}$, and $75^{th}$ percentiles of $l_{p}$ and $m_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; interquartile range of $l_{p}$ and $m_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; persistent entropy of $l_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$; and ALPS statistic of $l_{p}$ for $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$.
@@
4 Simulation studies
To demonstrate the utility of our method, we consider a simple simulated setup and evaluate our method against other well-known methods in the multivariate changepoint literature. We consider both a straightforward and a more difficult (noisier) changepoint problem for a topological change in a rather short image series. This is to show that our topological feature embedding requires relatively little data to perform well, and that our changepoint method also performs well and is robust in this “small” data setting. The following two subsections discuss the ability of our method to detect complicated changes in multivariate data of mixed type as well as changes in covariance.
4.1 Detecting a topological change in image series